OpenStudy (anonymous):

Oh goodness, I have another question. Solve the equation and give the highest value. (3x+2) (x-5)=0 each time I try to solve it I end up with something funky like 3x^2-13x=10, which I don't think is right. Can someone tell me where I'm going wrong and help me find the answer?

OpenStudy (agent0smith):

If a*b = 0, then either a=0 or b=0 (3x+2) (x-5)=0 either 3x+2 = 0 or x-5 = 0 Solve both.

OpenStudy (anonymous):

The highest value for y=0 is one of the two solutions to the binomial.

OpenStudy (anonymous):

So in the end the correct answer would be x=5? :) @agent0smith @larryboxaplenty

OpenStudy (agent0smith):

The "give the highest value" is a bit odd. Does it just mean the largest value of the two solutions? Then yes, x=5. Though, that's not something that'd normally be asked, as opposed to just "solve the equation"

OpenStudy (anonymous):

The "highest" value overall if infinity, due to the domain/range of the equation

OpenStudy (agent0smith):

It's not a function, though, it's an equation equal to zero, so it doesn't have a domain or range. There's no highest value, unless they mean just the larger value of the two solutions, which is an odd thing to ask for since it's kinda meaningless.